On the same track, Train A departs southbound traveling at 50km/h, while Train B departs northbound traveling at the same speed. They begin at a distance of 10km apart. At the time of departure, a monarch butterfly starts to fly from Train A to Train B at a speed of 100km/h. When the butterfly reaches Train B, it turns around and flies towards Train A. It continues this until the Trains collide. What is the total distance the butterfly has fluttered its wings before the trains reach their demise?
Answer by stanbon(75887)

On the same track, Train A departs southbound traveling at 50km/h, while Train B departs northbound traveling at the same speed. They begin at a distance of 10km apart. At the time of departure, a monarch butterfly starts to fly from Train A to Train B at a speed of 100km/h. When the butterfly reaches Train B, it turns around and flies towards Train A. It continues this until the Trains collide. What is the total distance the butterfly has fluttered its wings before the trains reach their demise?
-----------------
Determine the time elapse till the trains meet.
Closing speed = (50+50) = 100 km/hr
Distance = 10 km
Time = dist/rate = 10/100 = 0.1 hr
---
Distance the butterfly flies = rate*time = (100 km/hr)*0.1 hr = 10 km
===========
Cheers,
Stan H.
==============